OpenStudy (anonymous):

The length of a rectangle is 4 more than the width. The area is 5 square inches. Find the dimensions. Please walk me through it, I want to find the answer, I just need help getting there. :)

OpenStudy (anonymous):

If the width is x, then the length is x+4. You have to multiply both of them together to get the area, 5. Can you write the equation for this?

OpenStudy (anonymous):

5=x(x+4)?

OpenStudy (anonymous):

Exactly! Can you expand that now?

OpenStudy (anonymous):

\[5=x^2+4x?\]

OpenStudy (anonymous):

Good, now you need to move the 5 onto the other side so you can factor and solve for x.

OpenStudy (anonymous):

\[x^2+4x-5=0\]

OpenStudy (anonymous):

Can you factor that?

OpenStudy (anonymous):

What are the factors of -5 that add up to 4?

OpenStudy (anonymous):

5,-1

OpenStudy (anonymous):

x+5=0 x-1=0 -5-5 +1 +1 x=-5 x=1

OpenStudy (anonymous):

Right, so it is't possible for lengths to be negative so the answer is x=1. If the width is 1, then the length would be?

OpenStudy (anonymous):

uhh, 5?